# Carve3D: Improving Multi-view Reconstruction Consistency for Diffusion Models with RL Finetuning

### 21 Dec 2023 v1

@@ Appendix E Ablation Studies
As shown in Figure 9, when the bounding box normalization is removed from MRC, the model would trivially increase the reward by reducing the size of the foreground object on the white background. This would lead to the model generating images containing only the white background, after longer finetuning. With bounding box normalization, the model would learn the harder task of improving the reconstruction consistency of the multiview images.
@@
As shown in Figure 10 Our KL divergence regularization does not sacrifice the model’s efficiency on improving its reward. Without KL divergence regularization, the KL divergence grows much faster. As discussed in Section 4.2, this leads to degraded object identity and loss of texture details.
Appendix F Limitations and Future Work
@@
Carve3D is limited by the reconstruction quality of Sparse View LRM [24, 17]. Because its reconstruction is not perfect, this leads to non-zero MRC metric on GT views as shown in Figures 14, 15 and 16. Due to this limitation of Sparse View LRM, Carve3D RL finetuned model can produce less high-frequency details than the base model in order to lower the image distance to the NeRF rendered views. This might be solved by using a future sparse view reconstructor that can preserve more details or training a dedicated model for computing MRC.
@@
Further increasing data size and batch size to improve generalization of the improved reconstruction consistency is possible. However, in this work, we are limited by the high computation cost of SDXL [39], Instant3D’s 100 denoising steps, and the high number of samples needed in DDPO. A few concurrent works could address this challenge. It is possible to substantially reduce the computation cost by switching to Consistency Models for one/few-step inference (e.g., LCM-LoRA [30]). In addition, we can also switch from DDPO to direct backpropagation of reward (e.g. Align-Prop [41], and DRaFT [11]) to reduce the number of samples needed. We leave these extensions as future work.
Figure 11: Qualitative comparison of MVDream, Instant3D with 10k, 20k, and 100k SFT steps, and Carve3D (five columns) on four prompts (four blocks separated by dotted line). In each block, we show their generated multi-view images in the 2-by-2 grid (top), reconstructed NeRF and extracted mesh (bottom) when given the prompt (middle). When compared to the base Instant3D-10K: Carve3D maintains the detailed texture and provides improved multi-view consistency and higher quality NeRF; in contrast, the models with prolonged SFT of 20K and 100K steps exhibit worse level of details and realism, while only providing slightly improved consistency.
@@
Figure 12: Qualitative comparison of Instant3D (the base model) and Carve3D (the model finetuned from Instant3D) on $12$ prompts (in 12 blocks separated by dotted line). In each block, we show the their generated multi-view images in the 2-by-2 grid (top), the reconstructed NeRF and the extracted mesh (bottom) when given the prompt (middle). We draw red boxes on the NeRF and the extracted mesh to highlight the artifacts in the NeRF and the mesh, resulting from the inconsistencies in the multi-view images. Carve3D maintains the detailed texture and provides improved multi-view consistency and higher quality NeRF than the base Instant3D.
@@
Figure 13: Diverse results from original Instant3D (left) and our Carve3D (right) on 4 prompts (in 4 blocks separated by the dotted line). In each block, we show the their generated multi-view images in the 2-by-2 grid (top), the reconstructed NeRF and the extracted mesh (bottom) when given the prompt (middle). Our RLFT does not compromise the diversity of the base Instant3D model, while improving the consistency.

### 09 Apr 2024 v2

@@ Appendix E Ablation Studies
As shown in Fig. 9, when the bounding box normalization is removed from MRC, the model would trivially increase the reward by reducing the size of the foreground object on the white background. This would lead to the model generating images containing only the white background, after longer finetuning. With bounding box normalization, the model would learn the harder task of improving the reconstruction consistency of the multi-view images.
@@
As shown in Fig. 10 Our KL divergence regularization does not sacrifice the model’s efficiency on improving its reward. Without KL divergence regularization, the KL divergence grows much faster. As discussed in Sec. 4.2, this leads to degraded object identity and loss of texture details.
Appendix F Broader Impacts and Future Work
Our Multi-view Reconstruction Consistency (MRC) metric can serve as a valuable tool for evaluating any multi-view generative methods and guiding future developments in the field. Although we only demonstrate our Reinforcement Learning Finetuning (RLFT) with MRC on one multi-view diffusion model [24], it can be directly adapted to other text-to-multi-view diffusion models; such adaptation only requires tuning a few hyperparameters related to the scaling laws for diffusion model RLFT (Sec. 4.2.3). Our surprising finding behind the choice of REINFORCE [54] over PPO [45] for better training stability could also be applied in broader RLFT scenarios.
As AI models grow more powerful, it becomes more important to evaluate and improve their safety and reduce their bias. RLFT has been widely used for Large Language Model (LLM) alignment as it allows models to be finetuned with hard-to-specify objectives and its results are generalizable without undermining the base model’s knowledge. As the first work to use RLFT for text-to-3D and on diffusion models at the SDXL scale, we hope Carve3D can inspire more alignment research in the computer vision community.
@@ Appendix F Broader Impacts and Future Work
Carve3D is limited by the reconstruction quality of Sparse View Large Reconstruction Model (LRM) [24, 17]. Because its reconstruction is not perfect, this leads to non-zero MRC metric on GT views as shown in Figs. 14, 15 and 16. Due to this limitation of Sparse View LRM, Carve3D RL finetuned model can produce less high-frequency details than the base model in order to lower the image distance to the NeRF rendered views. This might be solved by using a future sparse view reconstructor that can preserve more details or training a dedicated model for computing MRC.
@@
Further increasing data size and batch size to further improve reconstruction consistency is possible. However, in this work, we are limited by the high computation cost of SDXL [39], Instant3D’s 100 denoising steps, and the high number of samples needed in DDPO. A few concurrent works could address this challenge. It is possible to substantially reduce the computation cost by switching to Consistency Models for one/few-step inference (e.g., LCM-LoRA [30]). In addition, we can also switch from DDPO to direct backpropagation of reward (e.g. Align-Prop [41], and DRaFT [11]) to reduce the number of samples needed. We leave these extensions as future work.
Figure 11: Qualitative comparison of MVDream [46], Instant3D [24] with 10K (the base model), 20K, and 100K SFT steps, Carve3D (ours, finetuned from Instant3D-10K), Zero123++ [47], and SyncDreamer [28] (7 models in 7 columns) on 4 prompts (in 4 rows, numbered as 1-4, separated by dotted line). In each row, we show their generated multi-view images in the 2-by-2 grid (denoted as MV), reconstructed NeRF and extracted mesh (denoted as RM) when given the text prompt (denoted as TP). MVDream, Zero123++, and SyncDreamer generates inconsistent multi-view images and reconstruction artifacts (highlighted in red). For each result, we use the same randomly sampled initial noise for all models to ensure the comparison is fair. We let Zero123++ and SyncDreamer to use one of Carve3D’s output multi-view images as their input image conditioning. Instant3D-10K, -20K, and -100K and Carve3D demonstrates progressively better multi-view consistency and reconstruction quality. Instant3D-10K, Carve3D, Zero123++, and SyncDreamer exhibits the best texture details and realism, whereas Instant3D-20K and -100K with prolonged SFT steps compromise those qualities.
@@
Figure 12: Qualitative comparison of Instant3D (the base model) and Carve3D (ours, finetuned from Instant3D) on $12$ prompts (in 12 blocks, numbered as 1-12, separated by dotted line). In each block, we show the their generated multi-view images in the 2-by-2 grid (denoted as MV), the reconstructed NeRF and the extracted mesh (denoted as RM) when given the text prompt (denoted as TP). For each result, we use the same randomly sampled initial noise for all models to ensure the comparison is fair. We draw red boxes on the NeRF and the extracted mesh to highlight the artifacts in the NeRF and the mesh, resulting from the inconsistencies in the multi-view images. Carve3D maintains the detailed texture and provides improved multi-view consistency and higher quality NeRF than the base Instant3D.
@@
Figure 13: Qualitative comparison of Instant3D (the base model) and Carve3D (ours, finetuned from Instant3D) on 4 prompts (in 4 rows, numbered as 1-4, separated by the dotted line) demonstrating diversity. In each row, we show 3 results from each model, including the generated multi-view images in the 2-by-2 grid (denoted as MV), the reconstructed NeRF and the extracted mesh (denoted as bottom) when given the prompt (denoted as middle). For each result, we use the same randomly sampled initial noise for all models to ensure the comparison is fair. Our RLFT maintains the diversity of the base Instant3D model, while improving the consistency.